OpenStudy (anonymous):

Find the LCM of 6x^2+30x and 3x^2+18x+15

OpenStudy (anonymous):

factor both 6x^2+30x=6x(x+5) 3x^2+18x+15=3(x+1)(x+5) so all of the factors are: 6x(x+1)(x+5)

OpenStudy (anonymous):

my second question is how to find the LCM of 21x^2 + 105x and 3x^2 + 24x + 45......I am totally lost.